Mrs. Santos wants to give her nephew two sweets from a jar she keeps on her kitchen table. There are 11 sweets in this jar.

7 of the sweets are red

4 of the sweets are orange

Calculate the probability that Mrs. Santos will take a sweet of each colour if she takes two sweets randomly from the jar.


1
Expert's answer
2021-04-29T15:56:10-0400

We wish to calculate the probability of picking two sweets of each color. We assume that there was picking without replacement.

P(red, orange) =P(red and orange) or p(orange or red)

="(\frac{7}{11}*\frac {4}{10})+(\frac {4}{11}*\frac{7}{10})"

="\frac{28}{55}"